# Data-centric Prediction Explanation via Kernelized Stein Discrepancy

### v1

@@ 4.1 Qualitative Evaluation
Following the convention of previous contributions in the literature, we show the prediction explanation instances for all candidate approaches side by side. Figure 3 shows three test cases of the CIFAR10 classification task that cover different classification outcomes, including high-confident correct prediction, low-confident correct prediction, and low-confident incorrect prediction. For both correct prediction cases, we are confident that HD-Explain provides a better explanation than others in terms of visually matching test data points. E.g. brown frogs in Figure 3 (a) and deer on the grass in Figure 3 (b). In contrast, for the misclassified prediction case (Figure 3 (c)), we note the HD-Explain produces examples that do not even belong to the same class as the predicted one. It is a good outcome as the method is not confident in providing consistent explanations. RPS also shows such inconsistency, which aligns with its claim (Yeh et al., 2018). The other two baseline methods do not offer such properties and still produce explanations that match the predicted label well. It is hard to justify how those samples support such prediction visually (since no clear shared pattern is obvious to us). In addition, it is interesting to see that Influence Function and TracIn produce near identical explanations, reflecting their similarity in leveraging the perturbation of model parameters.
Figure 4 provides additional insights into Ovarian Cancer histopathology and Brain Tumor MRI datasets. HD-Explain again shows a better explanation for producing training samples that appear similar to the test samples (note, for the semantic similarity should be referred to a medical practitioner). E.g. respecting the scanning orientation of test points in MRI as shown in Figure 4 (b). We note all baseline approaches tend to produce similar explanations to test samples belonging to the same classes. Rather than providing individual prediction explanations, those approaches act closer to per-class interpreters that look for class prototypes. To verify this observation further, we conducted a quantitative evaluation, as described in the next section.
@@ 4.2 Quantitative Evaluation
Now, we aim to evaluate the performance of the candidate explanation method through quantitative analysis. As explainability is often subjective, we limit our experiments to training datasets where ground-truth explanation samples are available to obtain fair evaluation results. Specifically, given a training data sample $(\mathbf{x}_{i},y_{i})$, we generate a test point $\mathbf{x}_{t}$ by adopting two image data augmentation methods:
@@
  • Horizontal Flip: $\mathbf{x}_{t}=\text{flip}(\mathbf{x}_{i})$, where we flip images horizontally that do not compromise the semantic meaning of images for the three datasets.
@@
Since the data augmentation is guaranteed to maintain prediction consistency, the ideally best explanation for the generated test point is the original data point $\mathbf{x}_{i}$ itself, considering their similarity. Hence, the quantitative evaluation could be a sample retrieval evaluation where Hit Rate measures the probability of successful retrieval. To provide reliable statistics, we created $30$ augmented test points for each training data point ($>10,000$ data points) in each dataset, resulting in more than $300,000$ independent runs.
@@
Figure 5(a) shows the hit rate comparison among candidate methods on the three image classification datasets under Noise Injection data augmentation. It was surprising to observe that existing methods face significant difficulty in retrieving the idea explanatory sample ($\leq 10\%$), even with such a simple problem setup; only HD-Explain (and its variant) produces a reasonable successful rate ($>80\%$). This observation raised our curiosity to know the underlying reason. Figure 5(b) shows the following experiments where we compute Coverage of unique explanations. The metric, coverage, shows the ratio of explanation samples that are unique over many test points. The purpose is to understand how likely a user may receive instance-specific explanations if they have many test points to interpret. It turns out existing solutions produce only 10% - 50% coverage – many test points receive the same set of explanations, disregarding their unique characteristic. This result reflects our previous observation in Figure 4, where we saw multiple duplicated explanations (as shown with red corner tags). In fact, the explanations are often dominated by the class labels; data points predicted as the same class would receive a similar set of explanations. In contrast, HD-Explain shows substantially higher coverage, providing a significant advantage for instance-level prediction explanation.
@@
Regarding computation efficiency, while we have summarized the scalability limitation of the candidate methods in Table 1, there was no computational efficiency evaluation conducted in previous works. To complement this lack, we recorded the wall clock execution time for previous quantitative experiments as shown in Figure 5(c). As expected, the Influence Function takes longer to return its explanation than other methods. HD-Explain and TracIn* are comparable. However, it is worth knowing that HD-Explain works on the whole model prediction, whereas TracIn* is configured only to compute the gradient of the predictive layer (due to the scalability issue of memory usage). Indeed, if we look at HD-Explain*, we note its computation time is much shorter than TracIn*. RPS shows a significant advantage in computation efficiency as it is 1) limited to the last layer of the neural network and 2) no auto-differentiation is needed for computing the training data influence.
@@
(a) Hit Rate (Noise Injection)
(b) Coverage (both)
(c) Hit Rate (Horizontal Flip)
(d) Execution Time (both)
@@ 4.3 Kernel Options
Our previous experiments used the Radial Basis Function (RBF) kernel by default. One may suggest another choice over kernels that better fit a particular application domain. In this experiment, we compare three well-known kernels on the three image classification datasets, including Linear, RBF, and Inverse Multi-Quadric (IMQ) kernels.
Figure 7 shows our experiment results under both data augmentation scenarios. We note that HD-Explain equipped with an IMQ kernel generally performs better than the RBF kernel regarding explanation quality (Hit Rate). The advantage is significant when the data augmentation scenario is Horizontal Flip (Figure 7 (c)), which appears more challenging than Noise Injection. In addition, the Coverage of explanation (Figure 7 (b)) aligns with the hit rate, where IMQ leads to a higher value. However, HD-Explain equipped with linear kernel runs much faster than the other options, as shown in Figure 7 (d), providing it with higher efficiency on large datasets. Indeed, if we cross-check the results with Figure 5, we note that even a Linear kernel is sufficient for HD-Explain to stand out from other baseline models considering its performance and efficiency. RBF and IMQ kernels often need similar computation time.
@@ 5 Conclusion
This paper presents HD-Explain, a Kernel Stein Discrepancy-driven neural classifier explanation method. The method is straightforward but powerful in terms of prediction explanation and data debugging tasks. The computation is efficient and can work with large machine learning models (e.g. ResNet) as a whole model explainer. Considering that many applications adopt pre-trained foundation models, we also introduced a variant of HD-Explain (HD-Explain*) that can work on the predictive layers such that the explanation would be limited to the model parts of interest. With empirical evidence demonstrated in this paper, we are confident that the method is practically effective for improving the transparency of Machine Learning predictive systems.
Impact Statements
This paper presents a prediction explanation method for neural network-based classifiers. While the ultimate goal of prediction explanation (and this work) is to improve the model transparency and trustworthiness, we note it could be misused to cause intentional prediction bias (e.g. removing high-influential training data points for specific purposes).
Appendix A Problem Definition Recap
@@
Appendix B Additional Derivation of Kernelized Stein Discrepancy
@@
$\mathbb{E}_{x\sim p}[\phi(x)\nabla_{x}\log p(x)+\nabla_{x}\phi(x)]=\mathbb{E}_{x\sim p}[\mathcal{A}_{p}\phi(x)]=0,\quad\forall f.$
@@
$\max_{\phi\in\mathcal{F}}\mathbb{E}_{x\sim{\color[rgb]{1,0.6,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6,0.2}\pgfsys@color@rgb@stroke{1}{0.6}{0.2}\pgfsys@color@rgb@fill{1}{0.6}{0.2}q}}[\mathcal{A}_{p}\phi(x)]=\max_{\phi\in\mathcal{F}}\mathbb{E}_{x\sim{\color[rgb]{1,0.6,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6,0.2}\pgfsys@color@rgb@stroke{1}{0.6}{0.2}\pgfsys@color@rgb@fill{1}{0.6}{0.2}q}}[\mathcal{A}_{p}\phi(x)]-\underbrace{\mathbb{E}_{x\sim{\color[rgb]{1,0.6,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6,0.2}\pgfsys@color@rgb@stroke{1}{0.6}{0.2}\pgfsys@color@rgb@fill{1}{0.6}{0.2}q}}[\mathcal{A}_{\color[rgb]{1,0.6,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6,0.2}\pgfsys@color@rgb@stroke{1}{0.6}{0.2}\pgfsys@color@rgb@fill{1}{0.6}{0.2}q}\phi(x)]}_{=0}=\max_{\phi\in\mathcal{F}}\mathbb{E}_{x\sim{\color[rgb]{1,0.6,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6,0.2}\pgfsys@color@rgb@stroke{1}{0.6}{0.2}\pgfsys@color@rgb@fill{1}{0.6}{0.2}q}}[\underbrace{\phi(x)}_{\text{projection coefficients}}\underbrace{(\nabla_{x}\log p(x)-\nabla_{x}\log{\color[rgb]{1,0.6,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6,0.2}\pgfsys@color@rgb@stroke{1}{0.6}{0.2}\pgfsys@color@rgb@fill{1}{0.6}{0.2}q}(x))}_{\text{score function difference}}]$
@@
$\max_{\phi\in\mathcal{F}}\mathbb{E}_{x\sim{\color[rgb]{1,0.6,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6,0.2}\pgfsys@color@rgb@stroke{1}{0.6}{0.2}\pgfsys@color@rgb@fill{1}{0.6}{0.2}q}}[\mathcal{A}_{p}\phi(x)]=\max_{\phi\in\mathcal{F}}\mathbb{E}_{x\sim{\color[rgb]{1,0.6,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6,0.2}\pgfsys@color@rgb@stroke{1}{0.6}{0.2}\pgfsys@color@rgb@fill{1}{0.6}{0.2}q}}[\mathcal{A}_{p}\sum_{i}w_{i}\phi_{i}(x)]=\max_{\phi\in\mathcal{F}}\sum_{i}w_{i}\mathbb{E}_{x\sim{\color[rgb]{1,0.6,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6,0.2}\pgfsys@color@rgb@stroke{1}{0.6}{0.2}\pgfsys@color@rgb@fill{1}{0.6}{0.2}q}}[\mathcal{A}_{p}\phi_{i}(x)],$
@@
$\max_{\phi}\langle\phi(\cdot),\mathbb{E}_{x\sim{\color[rgb]{1,0.6,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6,0.2}\pgfsys@color@rgb@stroke{1}{0.6}{0.2}\pgfsys@color@rgb@fill{1}{0.6}{0.2}q}}[\mathcal{A}_{p}k(\cdot,x)]\rangle_{\mathcal{H}},\quad\textit{s.t.}||\phi||_{\mathcal{H}}\leq 1.$
The optimal $\phi$ is therefore a normalized version of $\mathbb{E}_{x\sim{\color[rgb]{1,0.6,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6,0.2}\pgfsys@color@rgb@stroke{1}{0.6}{0.2}\pgfsys@color@rgb@fill{1}{0.6}{0.2}q}}[\mathcal{A}_{p}k(\cdot,x)]$. Hence, KSD is defined as the optimal between the distribution $p$ and $q$ with the optimal solution of $\phi$
$\mathbb{S}(q,p)=\mathbb{E}_{x,x^{\prime}\sim{\color[rgb]{1,0.6,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6,0.2}\pgfsys@color@rgb@stroke{1}{0.6}{0.2}\pgfsys@color@rgb@fill{1}{0.6}{0.2}q}}[\kappa_{p}(x,x^{\prime})],\quad\text{where}\quad\kappa_{p}(x,x^{\prime})=\mathcal{A}_{p}^{x}\mathcal{A}_{p}^{x^{\prime}}k(x,x^{\prime}).$
Appendix C Discussion: Intuition on why HD-Explain works
@@
    $[\nabla_{\mathbf{x}_{a}}\log f_{\theta}(\mathbf{x}_{a})_{y_{a}}||f_{\theta}(\mathbf{x}_{a})]^{\top}[\nabla_{\mathbf{x}_{b}}\log f_{\theta}(\mathbf{x}_{b})_{y_{b}}||f_{\theta}(\mathbf{x}_{b})]=\underbrace{\nabla_{\mathbf{x}_{a}}\log f_{\theta}(\mathbf{x}_{a})_{y_{a}}^{\top}\nabla_{\mathbf{x}_{b}}\log f_{\theta}(\mathbf{x}_{b})_{y_{b}}}_{\text{similarity of scores (input gradients)}}+\underbrace{f_{\theta}(\mathbf{x}_{a})^{\top}f_{\theta}(\mathbf{x}_{b})}_{\text{similarity of predictions}},$
@@
Appendix D Dataset Details

### v2

@@ 4.1 Qualitative Evaluation
Figure 3 shows three test cases of the CIFAR10 classification task that cover different classification outcomes, including high-confident correct prediction, low-confident correct prediction, and low-confident incorrect prediction. For both correct prediction cases, we are confident that HD-Explain provides a better explanation than others in terms of visually matching test data points e.g. brown frogs in Figure 3 (a) and deer on the grass in Figure 3 (b). In contrast, for the misclassified prediction case (Figure 3 (c)), we note the HD-Explain produces an example that does not even belong to the same class as the predicted one. This reflects a low confidence in model’s prediction for the particular test example and highlight a potential error in the prediction. RPS also shows such inconsistency in explanation, which aligns with its claim (Yeh et al., 2018). The other two baseline methods do not offer such properties and still produce explanations that match the predicted label well. It is hard to justify how those training samples support such prediction visually (since no clear shared pattern is obvious to us). In addition, it is interesting to see that Influence Function and TracIn produce near identical explanations, reflecting their similarity in leveraging the perturbation of model parameters.
(a) High-Conf Correct Predict
(b) Low-Conf Correct Predict
(c) Low-Conf Incorrect Predict
Figure 3: Qualitative evaluation of various example-based explanation methods using CIFAR10. We show three scenarios where the target model makes 1) a highly-confident prediction that matches ground truth label, 2) a low-confident prediction that matches ground truth label, 3) low-confident prediction that does not match ground truth label (which is a bird). For each sub plot, we show top-3 influential training data points picked by the explanation methods for the test example.
(a) High-Conf Correct Predict
(b) High-Conf Correct Predict
(c) Low-Conf Correct Predict
Figure 4: Qualitative evaluation of various example-based explanation methods using SVHN. We show three scenarios where the target model makes 1-2) a highly-confident prediction that matches ground truth label, 2) a low-confident prediction that matches ground truth label. The model predictive accuracy is near 100%, such that there is no incorrect prediction observed during our random search. For each sub plot, we show top-3 influential training data points picked by the explanation methods for the test example.
Figure 4 provides additional insights on SVHN dataset. HD-Explain again shows a better explanation for producing training samples that appear similar to the test samples (note for the semantic similarity, these explanations should be referred to a medical practitioner). For instance, the explanation of HD-Explain follows the similar environmental context of test points as shown in Figure 4 (a-b). To verify this observation further, we conducted a quantitative evaluation as described in the next section.
The qualitative evaluation for OCH and MRI datasets are given in the appendix due to the page limit of the main paper. The overall observations remain consistent with CIFAR10 and SVHN.
@@ 4.2 Quantitative Evaluation
In order to perform quantitative evaluation, we limit our experiments to datasets where ground-truth explanation samples are available. Specifically, given a training data sample $(\mathbf{x}_{i},y_{i})$, we generate a test point $\mathbf{x}_{t}$ by adopting two image data augmentation methods:
@@
  • Horizontal Flip: $\mathbf{x}_{t}=\text{flip}(\mathbf{x}_{i})$, where we flip images horizontally that do not compromise the semantic meaning of images.
@@
We created $30$ augmented test points for each training data point ($>10,000$ data points) in each dataset, resulting in more than $300,000$ independent runs. Since the data augmentation is guaranteed to maintain prediction consistency, the ideally best explanation for the generated test point is the original data point $\mathbf{x}_{i}$ itself. Hence, the quantitative evaluation could be a sample retrieval evaluation where Hit Rate measures the probability of successful retrieval.
@@
Figure 5(a) shows the hit rate comparison among candidate methods on the four image classification datasets under Noise Injection data augmentation. The existing methods face significant difficulty in retrieving the ideal explanatory sample ($\leq 10\%$), even with such a simple problem setup; only HD-Explain (and its variant) produces a reasonable successful rate ($>80\%$). We further investigate the diversity of explanations across a testset using the Coverage metric. Here, diversity indirectly reflects the granularity of an explanation when accumulated over the test dataset. Figure 5(b) shows the Coverage score, the ratio of explanation samples that are unique over many test points. It turns out that existing solutions produce only 10% - 50% coverage – many test points receive the same set of explanations, disregarding their unique characteristic. This result is in line with our previous observation in Figure 8, where we saw multiple duplicated explanations (as shown with red corner tags). We further observed that the explanations of baselines are often dominated by the class labels; data points predicted as the same class would receive a similar set of explanations. In contrast, HD-Explain shows substantially higher coverage, generating explanations that considers the unique characteristics of each test point.
@@
Regarding computation efficiency, while we have summarized the scalability limitation of the candidate methods in Table 1, there was no computational efficiency evaluation conducted in previous works. We recorded the wall clock execution time of each experiments as shown in Figure 5(c). As expected, the Influence Function takes longer to return its explanation than other methods. HD-Explain*, TracIn* and RPS, all use the last layer to generate explanation. RPS showed the lowest compute time since it does not require auto-differentiation for computing the training data influence. HD-Explain* showed the second best compute time and is efficient than TracIn* [^2] TracIn* is configured only to compute the gradient of the prediction layer due to its high memory requirements. and IF. HD-Explain considers the whole model for explanation and its compute time is not directly comparable to others. However, it shows better efficiency than IF across all datasets and is better than TracIn* on CIFAR10.
@@
(a) Hit Rate
(Noise Injection)
(b) Coverage
(c) Hit Rate
(Horizontal Flip)
(d) Execution Time
@@ 4.3 Kernel Options
We use the Radial Basis Function (RBF) as our default choice of kernel. However, another kernel may better fit a particular application domain. In this experiment, we compare three well-known kernels i.e. Linear, RBF, and Inverse Multi-Quadric (IMQ) on the three image classification datasets.
Figure 7 presents the results under both data augmentation scenarios. Overall, the IMQ kernel performs better than the RBF kernel regarding explanation quality (Hit Rate). The advantage is significant when the data augmentation scenario is Horizontal Flip (Figure 7c, which appears more challenging than Noise Injection. IMQ also showed better performance on Coverage (Figure 7b). The linear kernel performs worse compared to other kernels. However, it is substantially efficient than the others, as shown in Figure 7d, highlighting its utility on large datasets. Compared to the baselines presented in Figure 5, we note that the Linear kernel is sufficient for HD-Explain to stand out from other methods in both performance and efficiency.
@@ 5 Conclusion
This paper presents HD-Explain, a Kernel Stein Discrepancy-driven example-based prediction explanation method. We performed comprehensive qualitative and quantitative evaluation comparing three baseline explanation methods using three datasets. The results demonstrated the efficacy of HD-Explain in generating explanations that are accurate and effective in terms of their granularity level. In addition, compared to other methods, HD-Explain is flexible to apply on any layer of interest and can be used to analyze the evolution of a prediction across layers. HD-Explain serves as an important contribution towards improving the transparency of machine learning models.
Appendix A Appendix / supplemental material
Appendix B Problem Definition Recap
@@
Appendix C Additional Derivation of Kernelized Stein Discrepancy
@@
$\mathbb{E}_{x\sim p}[\phi(x)\nabla_{x}\log p(x)+\nabla_{x}\phi(x)]=\mathbb{E}_{x\sim p}[\mathcal{A}_{p}\phi(x)]=0,\quad\forall\phi.$
@@
$\max_{\phi\in\mathcal{F}}\mathbb{E}_{x\sim{\color[rgb]{1,0.6,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6,0.2}q}}[\mathcal{A}_{p}\phi(x)]$ $=\max_{\phi\in\mathcal{F}}\mathbb{E}_{x\sim{\color[rgb]{1,0.6,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6,0.2}q}}[\mathcal{A}_{p}\phi(x)]-\underbrace{\mathbb{E}_{x\sim{\color[rgb]{1,0.6,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6,0.2}q}}[\mathcal{A}_{\color[rgb]{1,0.6,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6,0.2}q}\phi(x)]}_{=0}$
$=\max_{\phi\in\mathcal{F}}\mathbb{E}_{x\sim{\color[rgb]{1,0.6,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6,0.2}q}}[\underbrace{\phi(x)}_{\text{projection coefficients}}\underbrace{(\nabla_{x}\log p(x)-\nabla_{x}\log{\color[rgb]{1,0.6,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6,0.2}q}(x))}_{\text{score function difference}}]$
@@
$\max_{\phi\in\mathcal{F}}\mathbb{E}_{x\sim{\color[rgb]{1,0.6,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6,0.2}q}}[\mathcal{A}_{p}\phi(x)]=\max_{\phi\in\mathcal{F}}\mathbb{E}_{x\sim{\color[rgb]{1,0.6,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6,0.2}q}}[\mathcal{A}_{p}\sum_{i}w_{i}\phi_{i}(x)]=\max_{\phi\in\mathcal{F}}\sum_{i}w_{i}\mathbb{E}_{x\sim{\color[rgb]{1,0.6,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6,0.2}q}}[\mathcal{A}_{p}\phi_{i}(x)],$
@@
$\max_{\phi}\langle\phi(\cdot),\mathbb{E}_{x\sim{\color[rgb]{1,0.6,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6,0.2}q}}[\mathcal{A}_{p}k(\cdot,x)]\rangle_{\mathcal{H}},\quad\textit{s.t.}||\phi||_{\mathcal{H}}\leq 1.$
The optimal $\phi$ is therefore a normalized version of $\mathbb{E}_{x\sim{\color[rgb]{1,0.6,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6,0.2}q}}[\mathcal{A}_{p}k(\cdot,x)]$. Hence, KSD is defined as the optimal between the distribution $p$ and $q$ with the optimal solution of $\phi$
$\mathbb{S}(q,p)=\mathbb{E}_{x,x^{\prime}\sim{\color[rgb]{1,0.6,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6,0.2}q}}[\kappa_{p}(x,x^{\prime})],\quad\text{where}\quad\kappa_{p}(x,x^{\prime})=\mathcal{A}_{p}^{x}\mathcal{A}_{p}^{x^{\prime}}k(x,x^{\prime}).$
Appendix D Discussion: Intuition on why HD-Explain works
@@
    $[\nabla_{\mathbf{x}_{a}}\log f_{\theta}(\mathbf{x}_{a})_{y_{a}}||f_{\theta}(\mathbf{x}_{a})]^{\top}$ $[\nabla_{\mathbf{x}_{b}}\log f_{\theta}(\mathbf{x}_{b})_{y_{b}}||f_{\theta}(\mathbf{x}_{b})]$
    $=\underbrace{\nabla_{\mathbf{x}_{a}}\log f_{\theta}(\mathbf{x}_{a})_{y_{a}}^{\top}\nabla_{\mathbf{x}_{b}}\log f_{\theta}(\mathbf{x}_{b})_{y_{b}}}_{\text{similarity of scores (input gradients)}}+\underbrace{f_{\theta}(\mathbf{x}_{a})^{\top}f_{\theta}(\mathbf{x}_{b})}_{\text{similarity of predictions}},$
@@
Appendix E Discussion: Relaxations of KSD Estimation
In section 3.1, we introduced multiple relaxations such that KSD estimation can support predictive model (as conditional distribution model) with discrete labels. However, we want to clarify that the relaxations introduced are not generally applicable to other context (e.g., Goodness of Fit) given their selective conditions on the data input distribution. In fact, the data distribution $P$ that generated the datasets $\{(\mathbf{x}_{i},y_{i})\}_{i}^{n}$ could be complex given the potentially intractable marginal distribution of $P(\mathbf{x})$. Our solution avoids the modelling of such complexity by only limiting the KSD discrepancy computation between model distribution $P_{\theta}$ and the sampled data distribution $P_{D}$ instead of touch original distribution $P$.
In particular, in the data-centric prediction explanation context, we only aim to extract training data points that are similar to the test sample, such that all inputs in the framework are valid (e.g., images) rather than random continuous inputs sampled from an unknown distribution $P$ that might be inevitable in other contexts. To apply the similar relaxation to other context, a thorough theoretical proof is needed, which is out of the scope of this research.
It is worth noting that the both relaxations introduced in this paper has corresponding theoretically rigid solutions (Yang et al., 2018; Jitkrittum et al., 2020) with the cost of computational overhead. While they are much more elegant, for explanation prediction purpose, we may prefer faster approximations. How to better incorporate the theoretically rigid solution in a more efficient way will be in our future research.
Appendix F Discussion: Prediction Explain Quality on Healthcare Dataset
(a) Overian Cancer Histopathology (OCH)
(b) Brain Tumor MRI
Figure 8: Qualitative evaluation of example-based explanation methods on Overian Cancer histopathology and Brain Tumor MRI datasets. We show two test data points that are predicted to belong to the same class in each dataset. Red triangle in the top right corner of an image shows the duplicate explanations across test samples.
Figure 8 provides additional insights into Ovarian Cancer histopathology and Brain Tumor MRI datasets. HD-Explain again shows a better explanation for producing training samples that appear similar to the test samples (note for the semantic similarity, these explanations should be referred to a medical practitioner). For instance, the explanation of HD-Explain follows the scanning orientation of test points in MRI as shown in Figure 8 (b). We note all baseline approaches tend to produce similar explanations to test samples belonging to the same classes. Rather than providing individual prediction explanations, those approaches act closer to per-class interpreters that look for class prototypes. To verify this observation further, we conducted a quantitative evaluation as described in the next section.
There is potential concern on interpreting our experiments on the two healthcare datasets provided, where no domain experts’ evaluation contained in this work.
In the quantitative evaluation (Figures 5 and 6), we highlight that HD-Explain demonstrates better explanation performance in terms of retrieving original data points that were used for data augmentation-based retrieval tests. This study is objective, requiring no domain knowledge for result validation.
Regarding the qualitative evaluation (as part of the visualization in Figure 4), we agree that the explanation quality of healthcare datasets needs insights from domain experts, whereas machine learning research in general often lacks corresponding authority to justify specific domain’s results. This challenge extends across the entire model explanation literature. Our objective is to facilitate the general population’s understanding, even in the absence of domain expert evaluation. It is important to recap that the provided explanations by HD-Explain are visually similar to the test data points. However, their deeper pathological interpretation requires further investigation by healthcare practitioners, which we encourage for future studies. Furthermore, experimental evaluation of our approach on CIFAR demonstrates that the effectiveness of our method is not limited to medical datasets and can be easily applied across other domains. Overall, we believe that qualitative evaluation aims to enhance understanding of the model’s behavior rather than serving as a performance justification.
Appendix G Dataset Details